Solve for A, B and r in the following equations:
A+B=9250
A+Br=9000
A+Br^2=8775
I have tried to cancel off the A's but the values I am getting do not make sense.
Could you please help?
Thank you.

Answer by Alan3354(69443)

Solve for A, B and r in the following equations:
A+B=9250
A+Br=9000
A+Br^2=8775
================
A+B=9250
A+Br=9000
------------ Subtract
B - Br = 250
B*(1-r) = 250
B = 250/(1-r)
-----------
A+B=9250
A+Br^2=8775
-------------- Subtract
B - Br^2 = 475
B(1-r^2) = 475
(1-r)*(1-r^2) = 475
r^3 - r^2 - r - 474 = 0
r = 8.1895258 (by graphing)
-----------
B = 250/(1-r)
B = -34.7728
------------
A = 9284.7728